If {xn} does not converge to L, show that there exist  > 0 and a subsequence {xnk
}
of {xn} such that |xnk − L| ≥  for each k ∈ N.
1
Expert's answer
2020-05-18T19:54:04-0400

Suppose (xn) does not converge to L.

So "\exists\isin" > 0 such that for each integer N there is an integer n = n(N) ≥ N

|xn − L| ≥ 0.

For N = 1 we obtain n1 = n(1) ≥ 1

such that |xn1 − L| ≥ 0.

let N = nk + 1 to obtain nk+1 = n(N) ≥ nk + 1

such that |xnk+1 − L| ≥ 0.

SO, there is a subsequence (xnk for k = 1, 2, · · ·

such that |xnk − L| ≥ 0 ∀ k = 1, 2, ·